OpenStudy (anonymous):

According to this property, 3(r + 4) = 3r + 12

OpenStudy (anonymous):

Yes, that's correct. What do you not understand or are you wanting confirmation?

OpenStudy (anonymous):

Commutative Distributive Multiplicative Identity Multiplicative of Zero Associative

OpenStudy (skullpatrol):

The property is the distributive property of multiplication with respect to addition.

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):

\[ 3(r + 4) = 3r + 12 \] $$a(b+c)=ab+ac$$